OpenStudy (anonymous):

Determine the solution to f(x) = g(x) using the following system of equations: f(x) = 3x − 23 g(x) = −4.5x + 7

OpenStudy (anonymous):

Just solve for the system of equations. What do you think is the answer?.

OpenStudy (anonymous):

x = 4 so now all you have to do is solve for y

OpenStudy (anonymous):

they asked for x

OpenStudy (snuggielad):

Nixy, where did you get x=4?

OpenStudy (anonymous):

Well, you have your answer now lol I just gave it to you lol

OpenStudy (snuggielad):

Why don't you explain that Nixy? That way the next time she can get it herself, that's the sign of a true helper.

OpenStudy (anonymous):

Well, the book she/he uses for math should explain it :-) Here are a few good links that explains system of equations http://www.mathsisfun.com/algebra/systems-linear-equations.html https://www.khanacademy.org/math/algebra-basics/core-algebra-systems/core-algebra-systems-tutorial/v/solving-linear-systems-by-graphing

OpenStudy (decentnabeel):

3x − 23=−4.5x + 7 add 4.5x and 23 on both sides 3x+4.5x=7+23 multiple by 10 on both sides \[\[10\cdot \:3x+10\cdot \:4.5x=10\cdot \:7+10\cdot \:23\]\] 75x=300 x=4

OpenStudy (decentnabeel):

are you understand @Naenae1999